# Think Outside the Policy: In-Context Steered Policy Optimization

### v1

@@ 6.2 Ablation Study
| GRPO | 83.00 | 25.10 / 22.19 | 66.87 | 42.28 | 47.89 |
| ICPO | 86.00 | 26.77 / 24.06 | 69.61 | 42.65 | 49.82 |
| OpenR1-Math-220k | |||||
| GRPO | 83.60 | 28.44 / 22.50 | 66.72 | 40.81 | 48.41 |
| ICPO | 86.80 | 31.25 / 26.25 | 70.39 | 44.12 | 51.76 |
@@
As shown in Table 5, results verify that IEF is consistently beneficial across training datasets of different difficulty levels. On the simpler Skywork dataset, ICPO improves GRPO by +1.93 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.35 average points. This demonstrates that IEF not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
6.3 Analysis on Source of Improvements
To better understand where the improvements of ICPO originate, we analyze the training dynamics of Qwen3-8B in Figure 6.
We observe that introducing IEF into mixed-policy GRPO drives the policy to deviate further from the reference distribution than vanilla GRPO, reflected by clear rises in entropy and KL divergence, accompanied by longer responses (Figure 6(a)). Importantly, this deviation is not uncontrolled drift: expert-conditioned rollouts encourage the model to explore expert-aligned yet previously unreachable regions of the solution space. As shown by the sharp peak in Figure 6(b), several prompts initially classified as All-Pass (i.e., groups whose GRPO rollouts are entirely correct) under GRPO begin producing more diverse reasoning strategies and transition to the Some-Pass group—precisely aligned with the onset of increased entropy and KL-loss. This indicates that IEF promotes deeper exploration beyond the solutions GRPO already masters. Meanwhile, ICPO consistently yields fewer Zero-Pass prompts than GRPO throughout training, indicating that expert guidance helps the model solve cases that GRPO fails on, thereby improving overall performance.
These observations together highlight that ICPO improves both the coverage and quality of successful rollouts without harming performance on easier problems. Nonetheless, the training dynamics demonstrate that ICPO enhances controllable exploration toward expert regions, which ultimately yields higher reasoning performance.
7 Conclusion
We present ICPO, a unified RLVR framework that enhances reasoning without relying on external expert models. Leveraging the inherent ICL capability of LRMs, ICPO introduces mixed-policy GRPO with IEF, which constructs expert-conditioned rollouts from existing datasets, improving data utilization and expanding exploration beyond the current policy distribution. To ensure stable optimization, ICPO further integrates ERRS to eliminate noisy off-policy trajectories and adopts RS to facilitate a smooth transition from expert-guided imitation to autonomous optimization. Experiments show that ICPO consistently improves RL performance, highlighting its promise as a scalable and general post-training paradigm for LRMs.
Ethics Statement
Use of AI Assistants
We certify that any use of AI tools, including ChatGPT, was strictly limited to linguistic refinement such as improving grammar, clarity, and style. All substantive ideas, analyses, and arguments presented in this work originate from the authors or from properly cited prior research.
Appendix A Prompt Construction
A.1 Prompt and Respones Format
Here we provide the detailed prompt format used for RL training, 0-shot and few-shot inference.
Appendix B Evaluation Settings
We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al. (2024), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), and MATH-500 Hendrycks et al. (2021). For AIME24, AIME25, and AMC23, which have relatively small test sets, we report Avg@32, while for the other benchmarks we report Pass@1. To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al. (2018), GPQA-Diamond Rein et al. (2023), and MMLU-Pro Wang et al. (2024), with multiple-choice options shuffled to prevent contamination. All evaluations are conducted using the LIMO framework Ye et al. (2025). During inference, we follow Yan et al. (2025) and set the generation temperature to 0.6 with a maximum response length of 8192 tokens. For few-shot ICL evaluation, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set using 5 different random seeds, and report the average performance across them.

### v2

@@ E.2 Ablation on Difficulty Levels
| GRPO | 83.0 | 25.1 / 22.2 | 66.9 | 42.3 | 47.9 |
| ICPO | 86.0 | 26.8 / 24.1 | 69.6 | 42.6 | 49.8 |
@@
We additionally train on a simpler dataset, Skywork-OR1-RL-Data444https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al. (2025), to verify the generalization ability of our IEF under different reasoning conditions. This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison. As shown in Table 12, results verify that ICPO is consistently beneficial across training datasets of different difficulty levels. On the simpler Skywork dataset, ICPO improves GRPO by +1.9 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.4 average points. This demonstrates that ICPO not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
Prompt Groups.
| Variant | MATH | AIME24/25 | AMC | Mnrv. | Avg. |
| GRPO | 83.0 | 25.1 / 22.2 | 66.9 | 42.3 | 47.9 |
| +IEF on Some-Pass | 82.0 | 26.9 / 26.0 | 67.8 | 43.4 | 49.2 |
| +IEF on Non-Pass | 82.4 | 25.42 / 23.33 | 67.19 | 42.28 | 48.12 |
| ICPO | 86.0 | 26.8 / 24.1 | 69.6 | 42.7 | 49.8 |
Table 13 further analyzes the effectiveness of ICPO by grouping prompts according to their initial rollout success rates. We denote prompts with a success rate of 0.0 as Non-Pass, and those with success rates between 0.0 and 1.0 as Some-Pass. ICPO applies implicit expert forcing (IEF) across all prompt groups.
We observe that expert guidance consistently improves performance for both Some-Pass and Non-Pass prompts. Moreover, the impact of IEF differs across prompt types and exhibits complementary benefits. For Some-Pass prompts, IEF enriches solution diversity while remaining within expert-aligned regions, thereby mitigating premature policy specialization. For Non-Pass prompts, IEF provides reliable expert guidance that enables the policy to solve prompts that cannot be handled by vanilla GRPO.
Taken together, these results demonstrate that ICPO enhances exploration for Some-Pass cases while enabling effective learning on challenging Non-Pass prompts.
Appendix F Token Rank Shift Analysis.
To quantify how different training methods alter model behavior at a fine-grained level, we follow Lin et al. (2023) and adopt token-rank shift as a token-level diagnostic metric. The core idea is to measure how unlikely the generated tokens of an aligned model are under the base model distribution.
Given a generated trajectory produced by an aligned policy, we replay the same trajectory under the base model and examine, at each generation step, how the aligned token is ranked according to the base model’s conditional distribution. Let denote the prefix up to position , and let be the token generated by the aligned model at step . We compute the base model logits and define the base-rank of as:
A rank of indicates that the aligned token coincides with the base model’s top-1 prediction, while larger ranks indicate increasing deviation from the base policy’s high-probability region.
For interpretability, we further categorize token ranks into three regimes: (1) Unshifted: ; (2) Marginally shifted: ; (3) Shifted: , where is a small threshold (e.g., ).
We define the shifted ratio as the fraction of tokens whose base-rank exceeds , which serves as a proxy for how frequently the aligned policy samples tokens that would be unlikely under the base model distribution.
Due to the varying lengths of reasoning trajectories, analyzing token rank shift using absolute token positions can be misleading. Instead, we normalize token positions by trajectory length and compute statistics over length percentiles. In Figure 7, we discretize the normalized positions into 100 bins and aggregate token shift metrics within each bin. By aggregating these statistics across token positions, token rank shift enables a fine-grained analysis of when and to what extent the aligned policy departs from the base policy during generation.